# The ideal structure of C-algebras of étale groupoids with isotropy groups of local polynomial growth

### v3

@@ Introduction
The problem simplifies significantly when the groupoid $\mathcal{G}$ is principal and therefore $\operatorname{Stab}(\mathcal{G})^{\mathrm{prim}}$ can be identified with ${\mathcal{G}^{(0)}}$. As was shown already by Effros and Hahn [EH] for transformation groupoids, we then get a homeomorphism $(\mathcal{G}\backslash{\mathcal{G}^{(0)}})^{\sim}\cong\operatorname{Prim}C^{*}(\mathcal{G})$, where $\sim$ denotes the $T_{0}$-ization, or the Kolmogorov quotient, obtained by identifying points with identical closures; in other words, $(\mathcal{G}\backslash{\mathcal{G}^{(0)}})^{\sim}$ is the quasi-orbit space for the action $\mathcal{G}\curvearrowright{\mathcal{G}^{(0)}}$. The same is true when $\mathcal{G}$ is assumed to be only essentially principal, meaning that every closed invariant subset $X\subset{\mathcal{G}^{(0)}}$ has a dense subset of points with trivial isotropy groups, see [Rbook, SW] and Section 4.5 below. This is closely related to the fact that in this case the induction map (*Introduction) factors through the unit space ${\mathcal{G}^{(0)}}$. When $\mathcal{G}$ is not essentially principal, this property no longer holds and we are forced to work with the entire space $\operatorname{Stab}(\mathcal{G})^{\mathrm{prim}}$.
@@
In our previous paper [CN3] we proved that (**∗ ∗Introduction) is indeed a homeomorphism for all amenable second countable étale groupoids with abelian isotropy groups. This covered a fairly large class of C-algebras including higher rank graph C-algebras and crossed products $\Gamma\ltimes C_{0}(X)$ defined by amenable actions $\Gamma\curvearrowright X$ with abelian stabilizers. Although there are plenty of nontrivial examples of such actions, see Section 5.1, it is nevertheless clear that unless $\Gamma$ is already abelian, the assumption of commutativity of the stabilizers is quite restrictive. At the same time we are not aware of any results in the literature that would describe $\operatorname{Prim}(\Gamma\ltimes C_{0}(X))$ for some infinite noncommutative group $\Gamma$ and arbitrary actions $\Gamma\curvearrowright X$.
@@
This in particular gives a description of $\operatorname{Prim}(\Gamma\ltimes C_{0}(X))$ for arbitrary actions $\Gamma\curvearrowright X$ of countable groups of local polynomial growth on second countable locally compact spaces. The theorem also covers arbitrary amenable actions $\Gamma\curvearrowright X$ of some groups of exponential growth. For example, we can take free products of groups of polynomial growth or, more generally, certain relatively hyperbolic groups, see Section 5.1.
In fact, our results are a bit more general than formulated above. First, we allow a strictly larger class of isotropy groups than groups of local polynomial growth, but we do not have a purely algebraic characterization of this class, see Section 4.1. Second, our main technical results do not need amenability of the entire groupoid and second countability, but then they can be used to describe only a part of $\operatorname{Prim}C^{*}(\mathcal{G})$. Note also that the above formulation does not have a form suggested by (**∗ ∗Introduction) or [MR0409720]*Conjecture 2, see Section 4.4 for further discussion.
@@
The class of FC-hypercentral groups is, in fact, the largest class of isotropy groups for which one can have such a simple description of the Jacobson topology, see Section 4.2. The theorem implies that (an adaptation of) [MR0409720]*Conjecture 2 is true for étale groupoids with FC-hypercentral isotropy groups and covers all $\mathcal{G}$ such that $C^{*}(\mathcal{G})$ is a type I C-algebra, see again Section 4.2, in certain sense completing the Mackey program in the étale case.
@@ 3.1. Necessary conditions for weak containment
Since every $\pi_{y}$ is weakly equivalent to a direct sum of irreducible representations, essentially the same proof shows that (1) implies the following property that is formally stronger than (2): for every neighbourhood $U$ of $x$, every finite set $F\subset\Gamma$ and every $\varepsilon>0$, there exist $y\in Y$, $\gamma\in\mathcal{G}_{y}$, an irreducible unitary representation $\theta$ of $\mathcal{G}^{y}_{y}$ and a Hilbert–Schmidt operator $T\colon H_{\rho}\to H_{\theta}$ such that $\|T\|_{2}=1$, $r(\gamma)\in U$, $\theta\prec\pi_{y}$ and, for every $g\in F$, we have either $W_{g}\cap\mathcal{G}^{r(\gamma)}_{r(\gamma)}=\emptyset$, or $W_{g}\cap\mathcal{G}^{r(\gamma)}_{r(\gamma)}=\{h\}$ for some $h$ and $\|T\rho(g)-\theta^{\gamma}_{y}(h)T\|_{2}<\varepsilon$. Correspondingly, in (3) we may in addition require that $\nu$ is a state on $C^{*}_{\theta^{\gamma}_{y}}(\mathcal{G}^{r(\gamma)}_{r(\gamma)})$ for some irreducible $\theta\prec\pi_{y}$. This could be used to slightly simplify a couple of subsequent arguments at the expense of having to keep track of more representations in the proof of Theorem 3.1.
@@ 4.1. Main result
Since $G$ is finitely generated, we thus see that $G$ does not belong to the class $\mathcal{M}$. Since $G$ embeds into groups from Example 4.3, this shows that neither $\mathcal{M}$ nor $\mathcal{M}_{0}$ is closed under taking subgroups.
@@ 5.1. Transformation groupoids defined by unipotent flows
For example, if $\Gamma$ is a free group, then the stabilizers are free as well by the Kurosh subgroup theorem, hence, being also amenable, they are either trivial or isomorphic to $\mathbb{Z}$. In this case the primitive spectrum is described by Theorem 4.6 or already by [CN3]*Corollary 2.8. Note that every connected Lie group $G$ such that $G/\operatorname{Rad}G$ is noncompact contains discrete subgroups isomorphic to the free group $F_{2}$, since $\operatorname{PSL}_{2}(\mathbb{R})$ is a subquotient of $G$ (because every noncompact real semisimple Lie algebra contains a copy of $\mathfrak{sl}_{2}(\mathbb{R})$) and $\operatorname{PSL}_{2}(\mathbb{Z})$ has plenty of free subgroups.
More generally, we can take a hyperbolic group $\Gamma$, as then every amenable subgroup of $\Gamma$ is virtually cyclic [GlH]*Théorème 8.37. Even more generally, $\Gamma$ can be hyperbolic relative to a family of subgroups $\mathcal{H}$ that are FC-hypercentral or of local polynomial growth (see, e.g., [MR2922380] for the relevant definitions). Indeed, by the strong Tits alternative [MR1313451]*Theorems 2T in this case every amenable subgroup of $\Gamma$ is either virtually cyclic or conjugate to a subgroup of some $H\in\mathcal{H}$.
Another possibility is to take arbitrary discrete $\Gamma\subset G$ and nilpotent $U$. Then the stabilizers are nilpotent as well and the primitive spectrum is again described by Theorem 4.6. For example, we can take $G=\operatorname{SL}_{n}(\mathbb{R})$ and $U$ to be the group $U_{n}(\mathbb{R})$ of unipotent upper triangular matrices, that is, upper triangular matrices with ones on the diagonal.
For $n=2$ we have $U_{2}(\mathbb{R})\cong\mathbb{R}$, so in this case, for any discrete subgroup $\Gamma\subset\operatorname{SL}_{2}(\mathbb{R})$, the stabilizers of the action $\Gamma\curvearrowright G/U$ are again either trivial or isomorphic to $\mathbb{Z}$, and we can apply [CN3]*Corollary 2.8. Note that neither of our results is needed if the action $\Gamma\curvearrowright G/U$ happens to be minimal, as then $\Gamma\ltimes C_{0}(G/U)$ is simple by [MR1258035]. By a classical theorem of Hedlund [MR1545946], see also [MR0393339], this is in particular the case when $\Gamma\subset\operatorname{SL}_{2}(\mathbb{R})$ is a uniform lattice. We remind that a lattice is called uniform or cocompact, if $\Gamma\backslash G$ is compact.

### v4

@@ Introduction
The problem simplifies significantly when the groupoid $\mathcal{G}$ is principal and therefore $\operatorname{Stab}(\mathcal{G})^{\mathrm{prim}}$ can be identified with ${\mathcal{G}^{(0)}}$. As was shown already by Effros and Hahn [EH] for transformation groupoids (and, in fact, even earlier by Glimm [MR0146297], if we accept surjectivity of the induction map), we then get a homeomorphism $(\mathcal{G}\backslash{\mathcal{G}^{(0)}})^{\sim}\cong\operatorname{Prim}C^{*}(\mathcal{G})$, were $\sim$ denotes the $T_{0}$-ization, or the Kolmogorov quotient, obtained by identifying points with identical closures; in other words, $(\mathcal{G}\backslash{\mathcal{G}^{(0)}})^{\sim}$ is the quasi-orbit space for the action $\mathcal{G}\curvearrowright{\mathcal{G}^{(0)}}$. The same is true when $\mathcal{G}$ is assumed to be only essentially principal, meaning that every closed invariant subset $X\subset{\mathcal{G}^{(0)}}$ has a dense subset of points with trivial isotropy groups, see [Rbook, SW] and Section 4.5 below. This is closely related to the fact that in this case the induction map (*Introduction) factors through the unit space ${\mathcal{G}^{(0)}}$. When $\mathcal{G}$ is not essentially principal, this property no longer holds and we are forced to work with the entire space $\operatorname{Stab}(\mathcal{G})^{\mathrm{prim}}$.
@@
In our previous paper [CN3] we proved that (**∗ ∗Introduction) is indeed a homeomorphism for all amenable second countable étale groupoids with abelian isotropy groups. This covered a fairly large class of C-algebras including higher rank graph C-algebras and crossed products $\Gamma\ltimes C_{0}(X)$ defined by amenable actions $\Gamma\curvearrowright X$ with abelian stabilizers. Although there are plenty of nontrivial examples of such actions, it is nevertheless clear that unless $\Gamma$ is already abelian, the assumption of commutativity of the stabilizers is quite restrictive. At the same time we are not aware of any results in the literature that would describe $\operatorname{Prim}(\Gamma\ltimes C_{0}(X))$ for some infinite noncommutative group $\Gamma$ and arbitrary actions $\Gamma\curvearrowright X$.
@@
This in particular gives a description of $\operatorname{Prim}(\Gamma\ltimes C_{0}(X))$ for arbitrary actions $\Gamma\curvearrowright X$ of countable groups of local polynomial growth on second countable locally compact spaces. The theorem also covers arbitrary amenable actions $\Gamma\curvearrowright X$ of some groups of exponential growth. For example, we can take free products of groups of local polynomial growth or, more generally, groups that are relatively hyperbolic groups with respect to a family of subgroups of local polynomial growth, see Section 4.1.
In fact, our results are a bit more general than formulated above. First, we allow a strictly larger class of isotropy groups than groups of local polynomial growth, but we do not have a purely algebraic characterization of this class, see again Section 4.1. Second, our main technical results do not need amenability of the entire groupoid and second countability, but then they can be used to describe only a part of $\operatorname{Prim}C^{*}(\mathcal{G})$. Note also that Theorem A does not have a form suggested by (**∗ ∗Introduction) or [MR0409720]*Conjecture 2, see Section 4.4 for further discussion.
@@
The class of FC-hypercentral groups is, in fact, the largest class of isotropy groups for which one can have such a simple description of the Jacobson topology, see Section 4.2. Theorem B describes in particular $\operatorname{Prim}(\Gamma\ltimes C_{0}(X))$ for actions of FC-hypercentral groups and for amenable actions of groups that are relatively hyperbolic with respect to a family of FC-hypercentral subgroups. The class of such relatively hyperbolic groups includes notably all lattices in semisimple Lie groups of real rank one. The theorem also covers all $\mathcal{G}$ such that $C^{*}(\mathcal{G})$ is a type I C-algebra, in certain sense completing the Mackey program in the étale case. It also implies that (an adaptation of) [MR0409720]*Conjecture 2 is true for étale groupoids with FC-hypercentral isotropy groups.
@@ 3.1. Necessary conditions for weak containment
Since every $\pi_{y}$ is weakly equivalent to a direct sum of irreducible representations, essentially the same proof shows that (1) implies the following property that is formally stronger than (2): for every neighbourhood $U$ of $x$, every finite set $F\subset\Gamma$ and every $\varepsilon>0$, there exist $y\in Y$, $\gamma\in\mathcal{G}_{y}$, an irreducible unitary representation $\theta$ of $\mathcal{G}^{y}_{y}$ and a Hilbert–Schmidt operator $T\colon H_{\rho}\to H_{\theta}$ such that $\|T\|_{2}=1$, $r(\gamma)\in U$, $\theta\prec\pi_{y}$ and, for every $g\in F$, we have either $W_{g}\cap\mathcal{G}^{r(\gamma)}_{r(\gamma)}=\emptyset$, or $W_{g}\cap\mathcal{G}^{r(\gamma)}_{r(\gamma)}=\{h\}$ for some $h$ and $\|T\rho(g)-\theta^{\gamma}(h)T\|_{2}<\varepsilon$. Correspondingly, in (3) we may in addition require that $\nu$ is a state on $C^{*}_{\theta^{\gamma}}(\mathcal{G}^{r(\gamma)}_{r(\gamma)})$ for some irreducible $\theta\prec\pi_{y}$. This could be used to slightly simplify a couple of subsequent arguments at the expense of having to keep track of more representations in the proof of Theorem 3.1.
@@ 4.1. Main result
Since $G$ is finitely generated, we thus see that $G$ does not belong to the class $\mathcal{M}$. Since $G$ embeds into groups from Example 4.3, this shows that neither $\mathcal{M}$ nor $\mathcal{M}_{0}$ is closed under taking subgroups.  $\diamond$
Theorem 4.2 covers in particular all transformation groupoids $\mathcal{G}=\Gamma\ltimes X$ defined by actions $\Gamma\curvearrowright X$ of countable groups of local polynomial growth on second countable Hausdorff locally compact spaces. It also covers the second countable transformation groupoids defined by amenable actions of groups $\Gamma$ that are relatively hyperbolic with respect to a family $\mathcal{H}$ of subgroups of local polynomial growth (see, e.g., [MR2922380] for the relevant definitions). Indeed, the stabilizers of such actions are amenable, while by the strong Tits alternative [MR1313451]*Theorems 2T every amenable subgroup of $\Gamma$ is either virtually cyclic or conjugate to a subgroup of some $H\in\mathcal{H}$, hence it is of local polynomial growth. For example, $\Gamma$ can be the free product of two countable groups of local polynomial growth or the free product of a hyperbolic group with a countable group of local polynomial growth. In this case the claim about amenable subgroups follows already from the Kurosh subgroup theorem and the fact that any amenable subgroup of a hyperbolic group is virtually cyclic [GlH]*Théorème 8.37. Since the Kurosh subgroup theorem is true for free products of arbitrary collections of groups, we can in fact take as $\Gamma$ the free products of countable collections of countable groups that are either of local polynomial growth or hyperbolic.
@@ 4.2. Groupoids with FC-hypercentral isotropy
Similarly to our discussion at the end of Section 4.1, Theorem 4.6 covers in particular all second countable transformation groupoids $\mathcal{G}=\Gamma\ltimes X$ defined by actions $\Gamma\curvearrowright X$ of FC-hypercentral groups, as well as the second countable transformation groupoids defined by amenable actions of groups that are relatively hyperbolic with respect to a family of FC-hypercentral subgroups. The latter class includes all lattices in connected semisimple Lie groups of real rank one: such lattices are known to be relatively hyperbolic with respect to a family of virtually nilpotent subgroups (see, for example, [HH]*Theorem 10.1 and [MR2154349]*Theorem 0.1).
@@ 5.1. Transformation groupoids defined by unipotent flows
For this to happen we may impose additional conditions on $\Gamma$. For example, as we discussed in Sections 4.1 and 4.2, we may require $\Gamma$ to be relatively hyperbolic with respect to a family of subgroups that are either FC-hypercentral or of local polynomial growth. Such groups $\Gamma$ include free groups and lattices in semisimple Lie groups of real rank one. We note in passing that every connected Lie group $G$ such that $G/\operatorname{Rad}G$ is noncompact contains discrete free subgroups, since $\operatorname{PSL}_{2}(\mathbb{R})$ is a subquotient of $G$ (because every noncompact real semisimple Lie algebra contains a copy of $\mathfrak{sl}_{2}(\mathbb{R})$) and $\operatorname{PSL}_{2}(\mathbb{Z})$ has plenty of free subgroups.
Another possibility is to take arbitrary discrete $\Gamma\subset G$ and nilpotent $U$. Then the stabilizers are nilpotent as well and the primitive spectrum is described by Theorem 4.6. For example, we can take $G=\operatorname{SL}_{n}(\mathbb{R})$ and $U$ to be the group $U_{n}(\mathbb{R})$ of unipotent upper triangular matrices, that is, upper triangular matrices with ones on the diagonal.
For $n=2$ we have $U_{2}(\mathbb{R})\cong\mathbb{R}$, so in this case, for any discrete subgroup $\Gamma\subset\operatorname{SL}_{2}(\mathbb{R})$, the stabilizers of the action $\Gamma\curvearrowright G/U$ are either trivial or isomorphic to $\mathbb{Z}$, and we can apply [CN3]*Corollary 2.8. Note that neither of our results is needed if the action $\Gamma\curvearrowright G/U$ happens to be minimal, as then $\Gamma\ltimes C_{0}(G/U)$ is simple. By a classical theorem of Hedlund [MR1545946], see also [MR0393339], this is in particular the case when $\Gamma\subset\operatorname{SL}_{2}(\mathbb{R})$ is a uniform lattice. We remind that a lattice is called uniform or cocompact, if $\Gamma\backslash G$ is compact.